Two sides of a triangle are 3ft and 6ft longer then the third side. If the perimeter of the triangle is 36ft find the lenght of each side
Found 2 solutions by lwsshak3, Alan3354:
Answer by lwsshak3(11628)

Two sides of a triangle are 3ft and 6ft longer then the third side. If the perimeter of the triangle is 36ft find the lenght of each side
let x= length of3rd side
x+3=2nd side
x+6 =1st side
2(x+3+x+6)=36
4x+18=36
4x=18
x=4.5
x+3=7.5
3rd side=4.5 ft
1st side=7.5 ft
2nd side= 10.5 ft

Answer by Alan3354(69443)

x + x+3 + x+6 = 36
3x + 9 = 36
3x = 27
x = 9
===============
---> 9, 12 & 15
P = 36